### v1

# Dual Attention Residual U-Net for Accurate Brain Ultrasound Segmentation in IVH Detection

Abstract
Abstract—— Brain intraventricular hemorrhage (IVH) remains a critical global concern, underscoring the urgent need for early detection and accurate diagnosis to improve survival rates among premature infants. Recent advances in deep learning have shown promising potential for computer-aided detection (CAD) systems to address this challenge. Our proposed approach consists of two powerful attention modules: the novel convolutional block attention module (CBAM) and sparse attention layer (SAL), integrated into a Residual U-Net model. The CBAM enhances the model’s ability to capture spatial relationships within local features. The SAL consists of sparse and dense attention branches. The sparse attention filters out the negative impacts of low query-key matching scores for aggregating features, whereas the original dense one ensures sufficient information flow through the network for learning discriminative representations. To evaluate our approach, we performed experiments using a widely used dataset of brain ultrasound images and the obtained results demonstrate its capability to accurately detect brain anatomies. Our approach achieves state-of-the-art performance with a dice score of 89.04% and IoU of 81.84% in the Brain US dataset in segmenting the ventricle region, showing its potential to help with precise brain anatomy detection. By leveraging the power of deep learning and integrating innovative attention mechanisms, our study contributes to ongoing efforts to improve brain anatomy detection. The source code can be found here: https://github.com/DanYuan001/BrainImgSegment
Keywords: Medical Image Segmentation, Res-UNet, Sparse Attention, Brain Anatomy
1 Introduction
According to the World Health Organization, 15 million babies are born prematurely each year [28]. For premature infants, one of the most common brain injuries is intraventricular hemorrhage (IVH) [33]. These hemorrhages result in ventricle dilation, which can cause serious brain damage if not properly treated. Therefore, we must continue to monitor the ventricle dilation to detect the occurrence of IVH as early as possible. From the point of view of machine vision technology, the task is to accurately segment the ventricle from an ultrasound image to obtain the anatomical structures of the brain.
In recent years, due to the advancements in deep learning research fields, medical image analysis and CAD systems have undergone a revolutionary transformation. There are many CAD systems that aim to automate the process of tumor segmentation in ultrasound images, assisting radiologists in accurate diagnosis. For example, the popularity of convolutional neural networks (CNNs) has increased for automatic detection and segmentation of tumor detection in ultrasound images. These models can learn feature representations from image data, enabling them to capture complex patterns and discriminative features [6]. Further, the study of various issues in medical images with the method of machine vision has led to the emergence of excellent models in the past, among which the most famous one is the U-Net series, such as U-Net [16], U-Net++ [17], V-Net [18] U-Net3+ [5], Res-UNet [25], and KiU-Net [12]. The principle of these U-Net variant models utilizes an encoder-decoder structure, with the encoder capturing context from the input images, and the decoder generating a segmentation map through upsampling. These models have demonstrated very good performance on various public and private medical datasets. However, CNNs lack the ability to capture the correlation of global image features due to the inherent spatial local inductive bias. More precisely, each convolutional kernel only processes a local subset of pixels in entire images, and forces the network to focus on local patterns rather than global context.
In a variety of computer vision tasks, attention mechanisms have produced promising results. The Attention U-Net [8], for instance, introduces attention gates to emphasize relevant features and suppress irrelevant ones. Other attention mechanisms, such as squeeze-and-excitation (SE), channel attention (CA) and spatial attention (SA) mechanism have also been applied to improve feature representations in learning frameworks [10], [11]. While U-Net variants excel in capturing local context, integrating global contextual information has been discovered as a shortcut to improve segmentation accuracy. Studies such as [13], [14] and [15], [16] have introduced different models to capture global contextual information effectively.
However, the key problem of the full attention mechanism is that it computes not only in some informative regions , but most of the uninformative ones of an image or feature map. The modified sparse attention mechanism in this paper has the ability to filter out the negative impact between queries and keys, and it is still able to capture the long-range data correlation, that is, its receptive field range is not weaker than that of the full attention models.
In light of the existing literature, we propose a novel segmentation method for detecting brain anatomies in brain ultrasound images. Our approach applies two attention mechanisms, namely the convolutional block attention module (CBAM) [31] and the sparse attention layer (SAL), which not only effectively capture context-aware features, spatial relationships, and global contextual information, but also neglect the correlations from uninformative regions. The entire architecture of the proposed segmentation model is shown in Figure 1.
The contributions of this work are as follows:
  • We propose a new sparse attention layer (SAL), which is selectively manipulate the correlations between data tokens, and is integrated into the proposed segmentation model;
  • We integrate into the model the CBAM attention mechanism which captures context-aware features through channel attention and spatial attention. This integration improves the contextual information and spatial relationships within local features, leading to more robust and accurate representations.
  • The combination of CBAM and SAL significantly improves the performance of the model on both the Brain US dataset and a supplementary dataset (Nuclei). The entire model achieves the state-of-the-art performance.
Figure 1: Block diagram of the proposed model used for medical image segmentation. An input image with dimensions 128 × 128 × 1 undergoes feature extraction through the encoder, and the decoder then performs upsampling on the encoded features to predict a binary mask of size 128 × 128 × 1. The in-between connections of the encoder and the decoder are accompanied by the addition of CBAM and SAL to enhance the performance.
2 Related Work
IVH is a pressing global health issue, and researchers have been exploring various methods to improve early detection and precise diagnosis to improve survival rates for premature infants. Among these efforts, deep learning-based systems have shown great promise in addressing this challenge. This section presents a review of the relevant literature that focuses on deep learning-based segmentation methods for detecting anatomies such as tumors and ventricles in ultrasound images. Deep learning methods have shown impressive achievements in diverse medical imaging tasks, such as image segmentation, classification, and detection [3], [12], [24], [20], [21], [22], [23].
In the study [24], the authors propose a novel method, called the confidence-guided brain anatomy segmentation network (CBAS), where the segmentation and corresponding confidence maps are estimated at different scales. In addition, they also introduce a technique which guides CBAS to learn the weights based on the confidence measure about the estimate.
In another study [12], the authors propose using an over-complete convolutional architecture where they project the input image into a higher dimension such that they constrain the receptive field from increasing in the deep layers of the network in order to detect small brain anatomical structures.
Chen et al. introduced AAU-Net [29], which is a hybrid adaptive module, combining convolutional layers with varying kernel sizes, channel self-attention, and spatial self-attention blocks to replace the traditional convolution operation.
CBAM, introduced in [31], demonstrated significant potential in improving CNNs to focus on relevant image regions. By integrating both channel and spatial attention mechanisms, CBAM enhances the representational power of CNNs and increases performance in various computer vision tasks [32].
In [35], the introduced method employed a deep ResNet architecture with a CBAM attention module to extract more comprehensive and in-depth features from pathological images. Positional attention mechanisms have gained attention in the medical imaging domain due to their ability to capture spatial relationships and contextual information within local features. The authors of [36] utilize positional attention in a multiscale framework to identify anatomical structures in medical images.
The residual U-Net architecture, introduced in [9], is a variant of the traditional U-Net model. Incorporating residual connections allows for better information flow during training and helps mitigate the vanishing gradient problem, leading to improved convergence and performance. The authors of [4] presented the Medical Transformer, a deep learning model designed for brain anatomy segmentation in ultrasound imaging. The model modifies the Res-UNet encoder with the gated axial-attention mechanism on different scales of the input images. However, the limitation of this model is the hypothesis that the attention computation is sufficient just along both the vertical and horizontal directions to save the expensive computation. In addition, another limitation of the model is the absence of a channel attention filter, which may restrict its capacity to emphasize the most critical feature layers.
In summary, while deep learning-based methods have shown promise in anatomical structure segmentation, there is still a need to explore more advanced attention mechanisms to improve the accuracy and robustness of the models.
3 Methodology
The research showcases an overall framework that integrates both CBAM and SAL based on residual U-Net (Res-UNet) scheme. The Res-UNet consists of the encoder and decoder, which work in a U-shaped pipeline(in Figure 1), i.e., extract features from the original image and generate segmentation results.
3.1 Encoder
In order to extract hierarchical features from the image, the encoder utilizes convolutional layers with 3×3 filters and a stride of 1. Batch normalization and ReLU function are applied after each layer to maintain feature stability. To avoid the problem of gradient vanishing during training and retain essential information, skip connections are utilized. As for the downsampling, 2×2 stride convolutional layers are applied. Additionally, the attention layers, such as SAL and CBAM, enhance the encoder features before connecting them with the decoder features via skip connections. More information on the attention layers can be found in the subsequent section.
3.2 Decoder
The decoder plays an important role in the reverse process of the encoder. In addition, the decoder also needs to combine upsampled feature maps with those from the attention layers at each level, including both low-level and high-level features. This helps refine features with convolutional layers, followed by batch normalization and ReLU activation. As a result, a higher-dimensional representation of the spatial relationships is obtained. Moreover, the SAL layer is incorporated in series to the CBAM, capturing global dependencies and filtering out the uninformative features in the process of attention computation.
3.3 Feature Fusion and Attention
Following feature extraction via the encoder, we conduct adaptive feature enhancement using attention mechanisms. The CBAM attention module comprises a sequence of efficient channel attention (CA) and spatial attention (SA) components, as well as SAL. The network details are shown in Figure 2.
Channel and Spatial Attention
Given a feature map $F\in\mathbb{R}^{H\times W\times C}$ , where H, W and C is the height, width and number of channels, respectively. CA assigns weights to the channels in the feature map, enhancing specific channels that more contribute to improving model performance. F is successively processed by a one-dimensional channel attention module and a two-dimensional spatial attention module , this process can be formulated as follows
$\begin{split}F_{C}=\sigma(mlp(gap(F))+mlp(gmp(F)))\otimes F,\\ F_{S}=\sigma(K^{7\times 7}[AvgPool(F_{C});MaxPool(F_{C})]),\\ F_{CS}=F_{S}\otimes F_{C}.\end{split}$ (1)
where $\sigma$ represents sigmoid activation function, and $\otimes$ denotes element-wise multiplication. $K^{7\times 7}$ is a large kernel with dilation of 4. $F_{C}\in\mathbb{R}^{H\times W\times C}$ and $F_{S}\in\mathbb{R}^{H\times W\times C}$ represent the features after the channel and spatial selections, and $F_{CS}\in\mathbb{R}^{H\times W\times C}$ is the output after these two modules. These two modules effectively capture channel-wise and spatial-wise dependencies, thereby allowing the model to focus on the relevant features, and improve its performance in image segmentation.
Figure 2: An illustration of CBAM and SAL blocks. CBAM and SAL are applied to the input feature F. The addition of the outputs of CBAM and SAL is fed to the decoder.
Sparse Attention Layer
As the vanilla Transformers compute all tokens inside the feature map, it may calculate many irrelevant regions. Thus, it not only computes non informative areas, but also introduces redundant and irrelevant features, which degrades the model performance. In order to cope with this issue, a type of sparse attention [1] layer is proposed to filter out the features with negative impacts on the map, i.e., the low relevance between queries and keys. Among all activation functions, the ReLU function is considered because it clears to zero in the negative range. Meanwhile, considering the over-sparsity of attention [2], a dense attention layer (DAL) is also involved to retain the crucial feature information. Unlike the sparse layer, the dense layer uses the softmax activation function. The idea of using this dual-branch scheme is how to reduce the noisy features and redundant information while properly retaining the informative one as much as possible. To this end, the entire framework fuses the features from these two branches in an adaptive fashion, i.e., adaptively takes features from branches and propagates them through the network.
Similar to the VIT [3] partitioning process, the nonoverlapping image patches are generated with a fixed size of $M\times M$, resulting in a flattened representation $F^{i}\in\mathbb{R}^{M^{2}\times C}$, where $F^{i}$ means the i-th patch feature. Then the matrices of query Q, key K and value V are generated from F as follows
$Q=FW_{Q},\hskip 14.22636ptK=FW_{K},\hskip 14.22636ptV=FW_{V}.$ (2)
where the linear projection matrices $W_{Q}$, $W_{K}$ and $W_{V}\in\mathbb{R}^{C\times d}$ are learnable and shared among all the patches. The attention computation can be defined as follows
$A=Activation(\frac{QK^{T}}{\sqrt{d_{K}}}+B)V.$ (3)
where A denotes the estimated attention matrix, and B is the learnable relative position bias.
The regular dense attention layer (DAL) is adopted in most of the existing works by substituting the Activation by softmax as follows
$Att_{1}=softmax(\frac{QK^{T}}{\sqrt{d_{K}}}+B).$ (4)
Since not all query tokens are closely relevant to the corresponding ones in keys, the computation of all similarities is not necessary for image segmentation. Intuitively, designing a sparse attention layer to build useful interactions among tokens could improve feature aggregation. A ReLU-based layer seems to be a plausible solution for obtaining sparsity in attention. It filters out the similarities with negative relevance, and propagates forward the most useful information:
$Att_{2}=ReLU^{2}(\frac{QK^{T}}{\sqrt{d_{K}}}+B).$ (5)
Note that ReLU-based SAL causes information loss, a certain compensation technique is often demanded to relax sparsity, which defies the motivation of using SAL over DAL.
Simply applying SAL will impose over-sparsity on the pipeline, i.e., the learned feature representation contains insufficient information for the following process. Conversely, simply using softmax-based DAL will inadvertently introduce redundant information in irrelevant regions, posing an interference in image segmentation. Therefore, rather than preferring one paradigm over the other, we propose a two-branch adaptive self-attention mechanism.
Then, by fusing both of the branches, the further weighted attention score can be defined as follows
$F_{SAL}=(\sum_{i=1}^{2}(\omega_{i}Att_{i}))V.$ (6)
where $\omega_{i}\in\mathbb{R},i=1,2$ are two normalized weights for adaptively modulating two branches.
This design ensures a better trade-off between redundant information from irrelevant areas that can be filtered and informative features that can be leveraged. Finally, the sparse attention layer takes the output from CBAM, and its resulting attention maps are concatenated to the decoders.
3.4 Loss Function
Dice loss [37], binary cross-entropy (BCE) loss [34], and focal loss [30] are popular loss functions in image segmentation tasks. These loss functions help guide the training process by quantifying the similarity between the ground truth and the predicted masks.
The Dice loss is derived from the Dice coefficient, also known as the F1 score. It measures the overlap between the ground truth and the predicted masks, in order to maximize their similarity. The Dice loss is computed using Eq. 7.
$l_{Dice}=1-\frac{2\times TP}{2\times TP+FP+FN}.$ (7)
In Eq. 7, TP represents the number of true positive pixels, FP represents the number of false positive pixels, and FN represents the number of false negative pixels.
The BCE loss is another widely used loss function for image segmentation. It minimizes the dissimilarity between the predicted and ground truth masks. The BCE loss is as the following Eq. 8.
$\begin{split}l_{BCE}(I,\hat{I})=-\frac{1}{WH}(\sum_{i=0}^{W-1}\sum_{j=0}^{H-1}[I(i,j)log(\hat{I}(i,j))\\ +(1-I(i,j))log(1-\hat{I}(i,j))]).\end{split}$ (8)
where I represents the ground truth label , $\hat{\textbf{{I}}}$ represents the predicted probability of the foreground.
Focal loss is used to address class imbalance in segmentation tasks and provide more focus on difficult cases. It assigns higher weights to misclassified pixels, while reducing the impact of easy pixels during training. The Focal loss is computed as follows
$l_{Focal}=-\frac{1}{N}(\sum_{i=1}^{N}\alpha(1-p_{i})^{\gamma}log(p_{i})).$ (9)
In Eq. 9, $\alpha$ is a balancing parameter to control the contribution of each class, and $\gamma$ is the focusing parameter that modulates the rate at which the loss focuses on hard classification cases.
In training, we use a combination of the Dice loss, BCE loss, and Focal loss as shown in Eq. 10 to guide the optimization process. By minimizing this combined loss during training, our model learns to accurately segment the desired regions of interest.
$l=l_{Dice}+l_{BCE}+l_{Focal}.$ (10)
4 Experiments and Results
4.1 Dataset Description
We use the brain anatomy segmentation (Brain US) [24], [27] dataset to evaluate our model. Intraventricular hemorrhage (IVH) resulting in enlargement of the brain ventricles is one of the main causes of preterm brain injury. The main imaging modality used for the diagnosis of brain disorders in preterm newborns is cranial ultrasound because of its safety and cost-effectiveness. In addition, the absence of septum pellucidum is an important biomarker for the diagnosis of septo-optic dysplasia. Automatic segmentation of the brain ventricles and septum pellucidum from these US scans is essential for an accurate diagnosis and prognosis of these diseases. After obtaining institutional review board (IRB) approval, US scans were taken from 20 different premature neonates (age ¡ 1 year). The total number of images collected was 1629 with annotations out of which 1300 were allocated for training and 329 for testing. We resize the images to 128×128 for all our experiments.
4.2 Evaluation Metrics
The performance of our segmentation model is evaluated using commonly used metrics: dice score (F1) and intersection over union (IoU) score. These metrics provide quantitative measures of the model’s ability to accurately delineate regions of interest.
Intersection over Union
IoU is a measure that quantifies the overlap between the predicted binary segmentation mask and the ground truth mask. It is calculated as the ratio of the intersection area between the two masks to their union area and is defined as follows
$IoU=\frac{TP}{TP+FP+FN}.$ (11)
Dice Score
The Dice score, also referred to as the F1 score, integrates both recall and precision into a single value for evaluation and is defined as follows
$F_{1}=\frac{2\times TP}{2\times TP+FP+FN}.$ (12)
4.3 Experimental Setup
We have developed our segmentation model using Python and have used the TensorFlow and Keras libraries for implementation. To speed up training, we used the high performance NVIDIA RTX 4060Ti 16G.
4.4 Hyperparameter Setup
Our model is trained for 150 epochs for the Brain US dataset, where every epoch represents one complete pass through the entire datasets. For the image size, we resize all images to a uniform size of 128×128 pixels, which are input into the model for segmentation. During the training phase, we use the Adam optimizer with a learning rate of 0.0001 to optimize the model. This choice of optimizer allows us to efficiently update the model’s parameters and enhance convergence during training. To ensure a comprehensive assessment, we have divided the data into a split 70% (train), 10% (test) and 20% (validation). We have leveraged the training set to optimize the model’s parameters, fine-tune hyperparameters using the validation set, and gauge the model’s ability to generalize to new data by utilizing the testing set.
4.5 Ablation Study
Upgrading Model from Scratch
A series of experiments are conducted to refine our segmentation model and evaluate the impact of upgrading modifications. These experiments include:
  1. i.
    Baseline U-Net model as the initial benchmark.
  2. ii.
    Residual U-Net model with skip connections.
  3. iii.
    Residual U-Net model with CBAM applied to the skip connections.
  4. iv.
    Proposed Res-UNet model with the strengths of CBAM and SAL.
The results in Table 2 show the efficacy of each modification in the upgrading way. Each model has been trained using the linear combination of Dice, BCE, and Focal loss. The CBAM improves performance, while the SAL further enhances accuracy and segmentation quality. Both CBAM and SAL increase Dice and IoU by around 3 points with respect to the baseline. Furthermore, compared to the pure Res-UNet case, the dual attention modules between encoders and decoders contribute about 1.5 to 2 point improvement in terms of both metrics.
Figure 3: Results of the ablation study indicate the improvement in model performance with each experimental modification. GT and PM are Ground Truth and Predicted Mask, respectively. Heatmaps demonstrate the improvement of the model’s performance in focusing on the region of interest.The SAL module filters out the negative impact of query-key matching scores so that the heatmap looks monotonic in the uninformative regions. Note that there is no attention map for case i and case ii.
Figure 3 visually illustrates the enhancement achieved with attention mechanisms. The combination of CBAM and SAL results in improved performance for both large and small regions of interest, refining feature representations, and capturing both global and local spatial dependencies for accurate segmentation. Compared with the CBAM-only case (iii), the full model (iv) performs the segmentation more precisely because the SAL module filters out the ’disturbing’ information from the attention map. Furthermore, the heatmap visualization of the proposed model illustrates the spatial regions where it places more emphasis, showing a close resemblance to the ground truth regions for the Brain US dataset.
Through these experiments and analysis, we are able to improve our segmentation model iteratively, identifying the most effective modifications and attention mechanisms. These advances make a significant contribution to enhancing the accuracy and robustness of our model, positioning it as an advanced solution for segmenting ultrasound images.
Combinations of Loss Functions
We have performed a series of experiments utilizing different loss combinations and have consolidated the results in Table 1. This table highlights the ablation study we have conducted on the loss functions employed to train our model. According to the results in Table 1, we have determined the most efficient loss function during model training and selected the optimal model configuration. Our findings reveal that the combination of Dice, BCE, and focal loss yields the highest performance.
Figure 4: Results of the proposed segmentation model on images of the Brain dataset. All the results are obtained by running their respective code.
4.6 State-of-the-art Comparisons
We have performed a comparison between our proposed method and several baseline models along with standard segmentation models. The comparative results, which comprehensively evaluate two metrics, are presented in Table 3. The models compared in the table are well known in the image segmentation field, such as FCN [7], U-Net [16], UNet++ [17], Res-UNet [25], Axial Attention U-Net [19] and Medical Transformer [4]. Our proposed method demonstrates superior performance across these models (see Table 3) in terms of Dice score and IoU, indicating better overall segmentation accuracy. To provide specific performance details, our proposed model achieves a Dice score of 89.04, indicating a higher level of similarity between the ground truth and the predicted segmentation masks. In addition, the IoU metric, with a value of 81.84, indicates the strong capability of the model to accurately delineate regions of interest. Significantly, the proposed method achieves the highest Dice score and IoU out of all the models, suggesting that it excels in terms of segmentation accuracy and overlap with ground truth.
Figure 5: Illustration of some of the failure cases of our model. The enclosed regions are the misclassified segmented masks. GT and PM represent the Ground Truth and Predicted Mask, respectively.
Table 1: Performance metrics of the proposed model with different loss functions.
Loss Functions Dice IoU
BCE Loss 70.73 74.91
Dice Loss 72.94 66.21
Focal Loss 67.16 61.40
BCE Loss + Dice Loss 87.32 78.89
BCE Loss + Focal Loss 74.61 76.27
Dice Loss + Focal Loss 83.24 75.02
BCE Loss + Dice Loss + Focal Loss 89.04 81.84
Table 2: Dice scores of the segmentation models. All values are in %. Bold values indicate superior performance. The results are in x(+y) format, where x is the score and y is the improvement over the baseline.
Metrics (i)U-Net[17] (ii)Res-UNet[27] (iii)Res-UNet (w/CBAM only) (iv)Res-UNet (w/CBAM and SAL(Ours))
Dice 85.37 87.50(+2.13) 87.97(+2.60) 89.04(+3.67)
IoU 79.31 79.61(+0.30) 80.64(+1.33) 81.84(+2.53)
Table 3: Performance comparison with standard segmentation models. All values are in %. Bold values indicate superior performance.
Metrics FCN[7] U-Net[16] U-Net++[17] Res-UNet[25] Axial Attention U-Net[19] Medical Transformer[4] Proposed (Ours)
Dice 82.79 85.37 86.59 87.50 87.92 88.84 89.04
IoU 75.02 79.31 79.95 79.61 80.14 81.34 81.84
Overall, the evaluation results show the superior performance of our proposed model compared to the state-of-the-art models. This confirms the effectiveness and robustness of our model in achieving accurate and precise segmentation results, positioning it as a promising solution for various segmentation tasks. Figure 4 shows the segmentation results of our proposed model, demonstrating its ability to segment the regions of interest in ultrasound images accurately. Although the visual result differences between our model and MedT are somehow tiny in most of the masks, it still achieves the best segmentation results by far. This indicates that the model focuses its attention on relevant areas, contributing to its accurate segmentation performance.
Through these comprehensive evaluations and visualizations, our proposed model showcases its potential to significantly improve brain anatomy detection and diagnosis, bringing us closer to the goal of early detection and enhanced patient care.
4.7 Weakness Analysis
Our model has demonstrated excellent performance in various image segmentation tasks, as depicted in Table 3 and Figure 4. However, it is essential to highlight the failure cases in which non-anatomical regions are misclassified as anatomical ones and vice versa, indicating that the precision and recall are relatively lower. It is important to acknowledge the complexity of the dataset used for the evaluation, which presents challenges in achieving perfect segmentation results. Figure 5 illustrates the difficult cases which our model encounters, resulting in deviations from the ground truth segmentation. These challenges may arise from dataset complexity, variations in image quality, or the presence of ambiguous features that are hard to accurately delineate. Despite these challenges, our proposed model demonstrates a valuable contribution to the field of brain anatomy segmentation.
4.8 Experimentation on the Nuclei Dataset
To evaluate the effectiveness of our proposed method, we have conducted assessments on another dataset named Nuclei, a well-known collection containing microscopic images of nucleus-stained slides and the corresponding ground truth annotations by expert pathologists. The average resolution for Nuclei images and the corresponding ground truth masks is 2000×2000 pixels. For the evaluation on this dataset, we have maintained consistency using the same set of hyperparameters that are employed in the evaluation of the Brain US dataset. However, the images should be converted to grayscale and resized to the same size as those in the Brain US dataset before being fed to the learning network. The segmentation results of the proposed model on the Nuclei dataset are shown in Figure 6. The visual results show that the overall performance is worse than those for the Brain US dataset because the segmentation regions are numbers of tiny particle-like areas, which make it hard for the attention to focus.
Figure 6: Predicted mask visualization of the proposed model on the Nuclei dataset. GT and PM represent the Ground Truth and Predicted Mask, respectively. The enclosed regions are the misclassified segmented masks.
5 Conclusion and Future Research
The current study presents an innovative segmentation technique called the attention-aided Res-UNet model to detect brain anatomies in ultrasound images. Our model has delivered satisfactory outcomes, revealing its potential to improve detection and diagnosis. Nonetheless, it is crucial to recognize the weakness of our proposed method.
In the weakness analysis section, several potential areas for future research were identified. One particularly promising direction involves investigating more hierarchical architectures that integrate multiple forms of stronger attention modules within the model. Such an approach has the potential to improve performance and deepen our understanding of complex brain anatomy detection challenges using complementary insights. Another direction for future research is to enhance datasets through techniques such as data enhancements, which can increase their diversity and size, thereby enhancing generalization and robustness.
In addition, evaluating the robustness of the model could provide valuable insight. Performing segmentation on other types of medical images, as we did in Nuclei dataset, would be beneficial in assessing the model’s versatility and applicability in diverse medical imaging tasks.
6 Acknowledgments
This project is supported by the Scientific and Technological Research Program of ChongQing Municipal Education Commission (Grant NO. KJQN202402609).

### v2

@@
 Abstract
-Abstract—— Brain intraventricular hemorrhage (IVH) remains a critical global concern, underscoring the urgent need for early detection and accurate diagnosis to improve survival rates among premature infants.
-Recent advances in deep learning have shown promising potential for computer-aided detection (CAD) systems to address this challenge.
-Our proposed approach consists of two powerful attention modules: the novel convolutional block attention module (CBAM) and sparse attention layer (SAL), integrated into a Residual U-Net model.
-The CBAM enhances the model’s ability to capture spatial relationships within local features.
-The SAL consists of sparse and dense attention branches.
-The sparse attention filters out the negative impacts of low query-key matching scores for aggregating features, whereas the original dense one ensures sufficient information flow through the network for learning discriminative representations.
-To evaluate our approach, we performed experiments using a widely used dataset of brain ultrasound images and the obtained results demonstrate its capability to accurately detect brain anatomies.
-Our approach achieves state-of-the-art performance with a dice score of 89.04% and IoU of 81.84% in the Brain US dataset in segmenting the ventricle region, showing its potential to help with precise brain anatomy detection.
-By leveraging the power of deep learning and integrating innovative attention mechanisms, our study contributes to ongoing efforts to improve brain anatomy detection.
-The source code can be found here: https://github.com/DanYuan001/BrainImgSegment
+Abstract—— Intraventricular hemorrhage (IVH) is a severe neurological complication among premature infants, necessitating early and accurate detection from brain ultrasound (US) images to improve clinical outcomes.
+Although recent deep learning methods offer promise for computerized diagnosis, challenges remain in capturing both local spatial details and global contextual dependencies critical for segmenting brain anatomies.
+In this work, we propose an enhanced Residual U-Net architecture incorporating two complementary attention mechanisms: the Convolutional Block Attention Module (CBAM) and a hybrid attention layers (HAL) made up of the Sparse Attention Layer (SAL) and Dense Attention Layer (DAL).
+The CBAM improves the model’s ability to refine spatial and channel-wise features, while the HAL introduces a dual-branch design, sparse attention filters out low-confidence query-key pairs to suppress noise, and dense attention ensures comprehensive information propagation.
+Extensive experiments on the Brain US dataset demonstrate that our method achieves state-of-the-art segmentation performance, with a Dice score of 89.04% and IoU of 81.84% for ventricle region segmentation.
+These results highlight the effectiveness of integrating spatial refinement and attention sparsity for robust brain anatomy detection.
+Code is available at: https://github.com/DanYuan001/BrainImgSegment.
 Keywords: Medical Image Segmentation, Res-UNet, Sparse Attention, Brain Anatomy
 1 Introduction
-According to the World Health Organization, 15 million babies are born prematurely each year [28].
+According to the World Health Organization, 15 million babies are born prematurely each year [31, 39, 40].
 For premature infants, one of the most common brain injuries is intraventricular hemorrhage (IVH) [35].
@@
 From the point of view of machine vision technology, the task is to accurately segment the ventricle from an ultrasound image to obtain the anatomical structures of the brain.
-In recent years, due to the advancements in deep learning research fields, medical image analysis and CAD systems have undergone a revolutionary transformation.
-There are many CAD systems that aim to automate the process of tumor segmentation in ultrasound images, assisting radiologists in accurate diagnosis.
-For example, the popularity of convolutional neural networks (CNNs) has increased for automatic detection and segmentation of tumor detection in ultrasound images.
-These models can learn feature representations from image data, enabling them to capture complex patterns and discriminative features [6].
-Further, the study of various issues in medical images with the method of machine vision has led to the emergence of excellent models in the past, among which the most famous one is the U-Net series, such as U-Net [16], U-Net++ [17], V-Net [18] U-Net3+ [5], Res-UNet [25], and KiU-Net [12].
-The principle of these U-Net variant models utilizes an encoder-decoder structure, with the encoder capturing context from the input images, and the decoder generating a segmentation map through upsampling.
-These models have demonstrated very good performance on various public and private medical datasets.
+Due to advances in deep learning networks, medical image analysis has undergone a tremendous revolution.
+There are many systems that conduct the anatomy segmentation process in ultrasound images, assisting radiologists in an accurate diagnosis.
+For instance, the breakthrough of convolutional neural networks (CNNs) in image classification and recognition has increased for automatic segmentation and recognition of anatomy in ultrasound images.
+The features of the various image data can be extracted and segmentation can be performed in ultrasound images [4].
+In addition, the U-Net series emerged in a variety of case studies using computer vision approaches in the medical imaging field, such as U-Net [15], U-Net++ [16], V-Net [17] U-Net3+ [5], Res-UNet [28] and KIU-Net[25].
+These variant models of U-Net consist of a basic encoder-decoder structure, with the encoder extracting features from the input images and the decoder producing a segmentation mask through upsampling in each level.
+The U-Net series have shown excellent performance in various public and private medical datasets.
 However, CNNs lack the ability to capture the correlation of global image features due to the inherent spatial local inductive bias.
-More precisely, each convolutional kernel only processes a local subset of pixels in entire images, and forces the network to focus on local patterns rather than global context.
-In a variety of computer vision tasks, attention mechanisms have produced promising results.
-The Attention U-Net [8], for instance, introduces attention gates to emphasize relevant features and suppress irrelevant ones.
-Other attention mechanisms, such as squeeze-and-excitation (SE), channel attention (CA) and spatial attention (SA) mechanism have also been applied to improve feature representations in learning frameworks [10], [11].
-While U-Net variants excel in capturing local context, integrating global contextual information has been discovered as a shortcut to improve segmentation accuracy.
-Studies such as [13], [14] and [15], [16] have introduced different models to capture global contextual information effectively.
+More precisely, each convolutional kernel only processes a local subset of pixels in the entire image and makes the network concentrate on local patterns rather than global context.
+In various computer vision tasks, attention mechanisms have produced promising results.
+The attention framework [18] and [4], for instance, introduces attention gates to emphasize relevant features and suppress irrelevant ones.
+Other attention mechanisms, such as sparse attention [1], have been utilized to highlight positive correlations between image tokens and filter out negative ones with the cooperation of the dense layer and the sparse layer.
+Some more attention mechanisms, such as the channel attention (CA) and spatial attention (SA) mechanism [10] have also been introduced to better describe the features extracted from images in learning networks.
+As we have discussed that CNNs are only capable of describing local context and so is the U-Net series, capturing global features of images has been a challenge in order to improve the vision system performance.
+Studies such as multilevel pyramid structure [14], multiscale approach like [13], and [41] have introduced various image ”zoom-in” models to achieve the goal.
 However, the key problem of the full attention mechanism is that it computes not only in some informative regions , but most of the uninformative ones of an image or feature map.
 The modified sparse attention mechanism in this paper has the ability to filter out the negative impact between queries and keys, and it is still able to capture the long-range data correlation, that is, its receptive field range is not weaker than that of the full attention models.
-In light of the existing literature, we propose a novel segmentation method for detecting brain anatomies in brain ultrasound images.
-Our approach applies two attention mechanisms, namely the convolutional block attention module (CBAM) [31] and the sparse attention layer (SAL), which not only effectively capture context-aware features, spatial relationships, and global contextual information, but also neglect the correlations from uninformative regions.
-The entire architecture of the proposed segmentation model is shown in Figure 1.
-The contributions of this work are as follows:
-  • We propose a new sparse attention layer (SAL), which is selectively manipulate the correlations between data tokens, and is integrated into the proposed segmentation model;
-  • We integrate into the model the CBAM attention mechanism which captures context-aware features through channel attention and spatial attention.
-This integration improves the contextual information and spatial relationships within local features, leading to more robust and accurate representations.
-  • The combination of CBAM and SAL significantly improves the performance of the model on both the Brain US dataset and a supplementary dataset (Nuclei).
-The entire model achieves the state-of-the-art performance.
+In this paper, we propose a novel segmentation method for detecting brain anatomies in brain ultrasound images.
+The approach is a Residual U-Net framework consisting of two attention mechanisms, namely the convolutional block attention module (CBAM) [33] and the hybrid attention layer (HAL), which not only better describe channel-wise features and spatial-wise relationships, but also neglect the correlations from uninformative regions and highlight informative ones.
+The general framework of the proposed segmentation model is shown in Fig.
+1.
+The contributions of this work are as follows.
+  • We utilize a revised hybrid attention layer (HAL), which selectively manipulate the correlations between data tokens, and is integrated into the proposed segmentation model;
+  • We integrate into the model the CBAM attention mechanism which captures both the channel-wise features and spatial-wise relationships, making the feature representations more accurate and interpretable.
+  • The series connection of CBAM and HAL in the entire framework significantly improves the performance of the model on both the Brain US dataset and a supplementary dataset (Nuclei).
+The entire model achieves state-of-the-art performance.
 Figure 1: Block diagram of the proposed model used for medical image segmentation.
 An input image with dimensions 128 × 128 × 1 undergoes feature extraction through the encoder, and the decoder then performs upsampling on the encoded features to predict a binary mask of size 128 × 128 × 1.
-The in-between connections of the encoder and the decoder are accompanied by the addition of CBAM and SAL to enhance the performance.
+The in-between connections of the encoder and the decoder are accompanied by the addition of CBAM and HAL to enhance the performance.
 2 Related Work
-IVH is a pressing global health issue, and researchers have been exploring various methods to improve early detection and precise diagnosis to improve survival rates for premature infants.
-Among these efforts, deep learning-based systems have shown great promise in addressing this challenge.
-This section presents a review of the relevant literature that focuses on deep learning-based segmentation methods for detecting anatomies such as tumors and ventricles in ultrasound images.
-Deep learning methods have shown impressive achievements in diverse medical imaging tasks, such as image segmentation, classification, and detection [3], [12], [24], [20], [21], [22], [23].
-In the study [24], the authors propose a novel method, called the confidence-guided brain anatomy segmentation network (CBAS), where the segmentation and corresponding confidence maps are estimated at different scales.
-In addition, they also introduce a technique which guides CBAS to learn the weights based on the confidence measure about the estimate.
-In another study [12], the authors propose using an over-complete convolutional architecture where they project the input image into a higher dimension such that they constrain the receptive field from increasing in the deep layers of the network in order to detect small brain anatomical structures.
-Chen et al.
-introduced AAU-Net [29], which is a hybrid adaptive module, combining convolutional layers with varying kernel sizes, channel self-attention, and spatial self-attention blocks to replace the traditional convolution operation.
-CBAM, introduced in [31], demonstrated significant potential in improving CNNs to focus on relevant image regions.
-By integrating both channel and spatial attention mechanisms, CBAM enhances the representational power of CNNs and increases performance in various computer vision tasks [32].
-In [35], the introduced method employed a deep ResNet architecture with a CBAM attention module to extract more comprehensive and in-depth features from pathological images.
-Positional attention mechanisms have gained attention in the medical imaging domain due to their ability to capture spatial relationships and contextual information within local features.
-The authors of [36] utilize positional attention in a multiscale framework to identify anatomical structures in medical images.
-The residual U-Net architecture, introduced in [9], is a variant of the traditional U-Net model.
-Incorporating residual connections allows for better information flow during training and helps mitigate the vanishing gradient problem, leading to improved convergence and performance.
-The authors of [4] presented the Medical Transformer, a deep learning model designed for brain anatomy segmentation in ultrasound imaging.
-The model modifies the Res-UNet encoder with the gated axial-attention mechanism on different scales of the input images.
-However, the limitation of this model is the hypothesis that the attention computation is sufficient just along both the vertical and horizontal directions to save the expensive computation.
-In addition, another limitation of the model is the absence of a channel attention filter, which may restrict its capacity to emphasize the most critical feature layers.
-In summary, while deep learning-based methods have shown promise in anatomical structure segmentation, there is still a need to explore more advanced attention mechanisms to improve the accuracy and robustness of the models.
+Intraventricular hemorrhage (IVH) remains a critical global health concern, particularly affecting premature infants.
+To improve early alarming and diagnosis, numerous deep learning-based approaches have been explored in recent years.
+These methods have shown substantial progress in segmentation and classification tasks in a wide range of medical imaging modalities [27, 25, 32, 4].
+For instance, the Confidence-Guided Brain Anatomy Segmentation Network (CBAS) [27] estimates both segmentation masks and corresponding confidence maps on multiple scales, and leverages a confidence-weighted learning mechanism to enhance model reliability.
+Similarly, the work in [25] proposes an over-complete convolutional architecture to project input images into higher dimensions, effectively controlling receptive field growth and enhancing small-structure detection.
+Recent efforts [22, 37] have begun to explore modular attention frameworks tailored to domain-specific constraints.
+IMAGDressing-v1 [23] introduces controllable visual pipelines integrating multilevel attention to achieve generation of structural consistency.
+Similarly, motion-aware modeling frameworks, such as MCDM [38] emphasize the importance of combining spatial structure, motion priors, and context-level awareness.
+Based on these insights, our work investigates the integration of channel-wise, spatial, and class-aware attention into a unified segmentation architecture optimized for IVH-related brain ultrasound imagery.
+In parallel, hybrid attention mechanisms have gained popularity for improving model focus and generalization.
+AAU-Net [32] integrates adaptive convolutions with channel and spatial self-attention to replace standard convolution operations, showing improvements in breast ultrasound segmentation.
+The Convolutional Block Attention Module (CBAM) [33] has also proven effective in guiding networks to attend to discriminative features, and has been widely adopted in both general vision and medical imaging domains.
+More recently, positional and multiscale attention mechanisms have been used to further model spatial dependencies [43], while transformer-based networks [4] have incorporated gated axial attention into residual U-Net backbones for improved context modeling.
+Despite these advancements, many existing architectures still struggle with balancing local spatial cues and global semantic representations, especially in ultrasound-based anatomical segmentation.
 3 Methodology
-The research showcases an overall framework that integrates both CBAM and SAL based on residual U-Net (Res-UNet) scheme.
-The Res-UNet consists of the encoder and decoder, which work in a U-shaped pipeline(in Figure 1), i.e., extract features from the original image and generate segmentation results.
-3.1 Encoder
-In order to extract hierarchical features from the image, the encoder utilizes convolutional layers with 3×3 filters and a stride of 1.
-Batch normalization and ReLU function are applied after each layer to maintain feature stability.
-To avoid the problem of gradient vanishing during training and retain essential information, skip connections are utilized.
-As for the downsampling, 2×2 stride convolutional layers are applied.
-Additionally, the attention layers, such as SAL and CBAM, enhance the encoder features before connecting them with the decoder features via skip connections.
-More information on the attention layers can be found in the subsequent section.
-3.2 Decoder
-The decoder plays an important role in the reverse process of the encoder.
-In addition, the decoder also needs to combine upsampled feature maps with those from the attention layers at each level, including both low-level and high-level features.
-This helps refine features with convolutional layers, followed by batch normalization and ReLU activation.
-As a result, a higher-dimensional representation of the spatial relationships is obtained.
-Moreover, the SAL layer is incorporated in series to the CBAM, capturing global dependencies and filtering out the uninformative features in the process of attention computation.
+The research shows an overall framework that integrates both CBAM and HAL based on the residual U-Net (Res-UNet) scheme.
+The Res-UNet consists of the encoder and decoder, which work in a U-shaped pipeline (in Fig.
+1), i.e.
+extract features from the original image and generate segmentation results.
+3.1 Encoder of Residual U-Net
+The encoder has four layers with multiple 3×3 convolution kernels and a stride of 1, which are common parameter configurations in U-Net.
+In each layer, convolutions are followed by batch normalization and the ReLU activation function for the network.
+After that, skip connections are introduced for the compensation to avoid the problem of gradient vanishing.
+In most of the U-Net framework, a ratio of 2 is applied for downsampling across the levels.
+In addition, the serialized attention modules HAL and CBAM, are fed with features output from the encoders to ensure the feature enhancements before connecting to the decoders via skip connections.
+Such a feature-intercepting scheme has been widely used in many case studies using U-Net frameworks, providing a way to refine the features before reaching the decoders.
+3.2 Decoder of Residual U-Net
+The decoder acts as the reverse process of the encoder.
+It also needs to combine upsampled feature maps with those from the attention layers at each level, including both low-level and high-level features.
+In addition, the batch normalization and activation functions are indispensable, with the same configuration as the encoder.
+Moreover, the HAL layer is incorporated in series with the CBAM, capturing global dependencies and filtering out the uninformative features in the process of attention computation.
 3.3 Feature Fusion and Attention
-Following feature extraction via the encoder, we conduct adaptive feature enhancement using attention mechanisms.
-The CBAM attention module comprises a sequence of efficient channel attention (CA) and spatial attention (SA) components, as well as SAL.
-The network details are shown in Figure 2.
-Channel and Spatial Attention
-Given a feature map $F\in\mathbb{R}^{H\times W\times C}$ , where H, W and C is the height, width and number of channels, respectively.
-CA assigns weights to the channels in the feature map, enhancing specific channels that more contribute to improving model performance.
-F is successively processed by a one-dimensional channel attention module and a two-dimensional spatial attention module , this process can be formulated as follows
-$\begin{split}F_{C}=\sigma(mlp(gap(F))+mlp(gmp(F)))\otimes F,\\ F_{S}=\sigma(K^{7\times 7}[AvgPool(F_{C});MaxPool(F_{C})]),\\ F_{CS}=F_{S}\otimes F_{C}.\end{split}$ (1)
-where $\sigma$ represents sigmoid activation function, and $\otimes$ denotes element-wise multiplication.
-$K^{7\times 7}$ is a large kernel with dilation of 4.
+Following feature extraction via the encoder, we perform adaptive feature enhancement using attention mechanisms.
+The CBAM attention module comprises a sequence of efficient channel attention (CA) and spatial attention (SA) components, as well as HAL.
+The network details are shown in Fig.
+2.
+3.3.1 Channel and Spatial Attention
+Given a feature map $F\in\mathbb{R}^{H\times W\times C}$ , where H, W, and C is the height, width, and number of channels, respectively.
+CA assigns weights to the channels on the feature map, enhancing specific channels that more contribute to improving model performance.
+F is successively processed by a one-dimensional channel attention module and a two-dimensional spatial attention module, this process is similar to [33] and [26], and can be formulated as follows
+$\begin{split}F_{C}=\sigma(MLP(AP(F))+MLP(MP(F)))\otimes F,\\ F_{S}=\sigma(K^{7\times 7}[AP(F_{C});MP(F_{C})]),\\ F_{CS}=F_{S}\otimes F_{C}.\end{split}$ (1)
+where $\sigma$ represents sigmoid activation function, AP denotes average pooling, MP denotes max pooling and $\otimes$ denotes element-wise multiplication.
+$K^{7\times 7}$ is a large kernel with a dilation of 4.
 $F_{C}\in\mathbb{R}^{H\times W\times C}$ and $F_{S}\in\mathbb{R}^{H\times W\times C}$ represent the features after the channel and spatial selections, and $F_{CS}\in\mathbb{R}^{H\times W\times C}$ is the output after these two modules.
-These two modules effectively capture channel-wise and spatial-wise dependencies, thereby allowing the model to focus on the relevant features, and improve its performance in image segmentation.
-Figure 2: An illustration of CBAM and SAL blocks. CBAM and SAL are applied to the input feature F. The addition of the outputs of CBAM and SAL is fed to the decoder.
-Sparse Attention Layer
-As the vanilla Transformers compute all tokens inside the feature map, it may calculate many irrelevant regions.
-Thus, it not only computes non informative areas, but also introduces redundant and irrelevant features, which degrades the model performance.
-In order to cope with this issue, a type of sparse attention [1] layer is proposed to filter out the features with negative impacts on the map, i.e., the low relevance between queries and keys.
+The CA module captures channel-wise dependencies and SA is for the spatial ones.
+Figure 2: An illustration of CBAM and HAL blocks.
+CBAM and HAL are applied to the input feature F.
+The addition of the outputs of CBAM and HAL is fed to the decoder.
+Note that the module ”HAL” is composed of a sparse layer and a dense layer.
+3.3.2 Hybrid Attention Layer
+As the Transformers compute all tokens inside the feature map, it may compute in many irrelevant areas.
+Thus, it introduces irrelevant features, which reduces the accuracy and computational efficiency of the model.
+In order to deal with this issue, a type of sparse attention [1] layer is proposed to filter out the features with negative impacts on the map, i.e., the low relevance between queries and keys.
 Among all activation functions, the ReLU function is considered because it clears to zero in the negative range.
-Meanwhile, considering the over-sparsity of attention [2], a dense attention layer (DAL) is also involved to retain the crucial feature information.
+In the meantime, there should be a dense attention layer (DAL) as a compensator if an oversparsity of attention occurs[2].
+The DAL retains crucial information about the features.
 Unlike the sparse layer, the dense layer uses the softmax activation function.
-The idea of using this dual-branch scheme is how to reduce the noisy features and redundant information while properly retaining the informative one as much as possible.
-To this end, the entire framework fuses the features from these two branches in an adaptive fashion, i.e., adaptively takes features from branches and propagates them through the network.
-Similar to the VIT [3] partitioning process, the nonoverlapping image patches are generated with a fixed size of $M\times M$, resulting in a flattened representation $F^{i}\in\mathbb{R}^{M^{2}\times C}$, where $F^{i}$ means the i-th patch feature.
-Then the matrices of query Q, key K and value V are generated from F as follows
-$Q=FW_{Q},\hskip 14.22636ptK=FW_{K},\hskip 14.22636ptV=FW_{V}.$ (2)
-where the linear projection matrices $W_{Q}$, $W_{K}$ and $W_{V}\in\mathbb{R}^{C\times d}$ are learnable and shared among all the patches. The attention computation can be defined as follows
+The idea of using this dual-branch scheme is to reduce the noisy information while retaining the informative one as much as possible.
+Therefore, the entire network copes with the fusion by using the features of these two branches and feeds them forward into successive modules.
+Similar to the VIT [3] partitioning process, the non-overlapping image patches are generated with a fixed size of $M\times M$, resulting in a flattened representation $F^{i}\in\mathbb{R}^{M^{2}\times C}$, where $F^{i}$ means the i-th patch feature.
+Then the matrices of query Q, key K, and value V are generated from F by multiplying it with linear projection matrices, respectively.
+Linear projection matrices $W_{Q}$, $W_{K}$, and $W_{V}\in\mathbb{R}^{C\times d}$ are learnable and shared among all patches.
+The attention can be computed as follows
 $A=Activation(\frac{QK^{T}}{\sqrt{d_{K}}}+B)V.$ (2)
-where A denotes the estimated attention matrix, and B is the learnable relative position bias.
-The regular dense attention layer (DAL) is adopted in most of the existing works by substituting the Activation by softmax as follows
+where A denotes the estimated attention result matrix, and B is the learnable relative position bias.
+The regular dense attention layer (DAL) is used in most of networks by substituting the activation by softmax as follows
 $Att_{1}=softmax(\frac{QK^{T}}{\sqrt{d_{K}}}+B).$ (3)
-Since not all query tokens are closely relevant to the corresponding ones in keys, the computation of all similarities is not necessary for image segmentation.
-Intuitively, designing a sparse attention layer to build useful interactions among tokens could improve feature aggregation.
-A ReLU-based layer seems to be a plausible solution for obtaining sparsity in attention.
-It filters out the similarities with negative relevance, and propagates forward the most useful information:
+Since any query token is not likely to be closely relevant to all keys, calculating all similarities is not necessary for image segmentation.
+Therefore, designing a sparse attention layer to build useful correlations among tokens could improve feature fusion.
+Replacement of activation with a ReLU-based layer appears to be a plausible way to obtain the sparsity of attention [1] [36].
+It feeds forward the useful information after filtering out the similarities with negative relevance:
 $Att_{2}=ReLU^{2}(\frac{QK^{T}}{\sqrt{d_{K}}}+B).$ (4)
-Note that ReLU-based SAL causes information loss, a certain compensation technique is often demanded to relax sparsity, which defies the motivation of using SAL over DAL.
-Simply applying SAL will impose over-sparsity on the pipeline, i.e., the learned feature representation contains insufficient information for the following process.
-Conversely, simply using softmax-based DAL will inadvertently introduce redundant information in irrelevant regions, posing an interference in image segmentation.
-Therefore, rather than preferring one paradigm over the other, we propose a two-branch adaptive self-attention mechanism.
-Then, by fusing both of the branches, the further weighted attention score can be defined as follows
-$F_{SAL}=(\sum_{i=1}^{2}(\omega_{i}Att_{i}))V.$ (6)
-where $\omega_{i}\in\mathbb{R},i=1,2$ are two normalized weights for adaptively modulating two branches.
-This design ensures a better trade-off between redundant information from irrelevant areas that can be filtered and informative features that can be leveraged.
-Finally, the sparse attention layer takes the output from CBAM, and its resulting attention maps are concatenated to the decoders.
-3.4 Loss Function
-Dice loss [37], binary cross-entropy (BCE) loss [34], and focal loss [30] are popular loss functions in image segmentation tasks.
-These loss functions help guide the training process by quantifying the similarity between the ground truth and the predicted masks.
-The Dice loss is derived from the Dice coefficient, also known as the F1 score.
-It measures the overlap between the ground truth and the predicted masks, in order to maximize their similarity.
-The Dice loss is computed using Eq.
-7.
-$l_{Dice}=1-\frac{2\times TP}{2\times TP+FP+FN}.$ (7)
-In Eq. 7, TP represents the number of true positive pixels, FP represents the number of false positive pixels, and FN represents the number of false negative pixels.
-The BCE loss is another widely used loss function for image segmentation. It minimizes the dissimilarity between the predicted and ground truth masks. The BCE loss is as the following Eq. 8.
+Note that a certain compensation technique is often demanded to overcome the sparsity. The feature fusion by building dual branches (SAL and DAL) is necessary.
+SAL, as an irrelevance filter, may cause insufficient information problems for the following process.
+In contrast, DAL inevitably introduces redundant information in many irrelevant regions, posing an interference in image segmentation.
+Therefore, we prefer a dual branch hybrid self-attention mechanism.
+Thus, by fusing both of the branches, the further weighted attention score can be defined as follows
+$F_{HAL}=(\sum_{i=1}^{2}(\omega_{i}Att_{i}))V.$ (5)
+where $\omega_{i}\in\mathbb{R},i=1,2$ are two normalized weights for modulating two branches.
+Finally, the attention layer takes the output from the CBAM, and its resulting attention maps are concatenated to the decoders.
+3.4 Loss Function Combinations
+There are quite a few loss functions suitable for image segmentation tasks.
+Dice loss, binary cross-entropy (BCE) loss, and focal loss are popular for our specific tasks.
+Although these loss functions are common in most textbooks, we still study the loss function selection strategy in the work [26], where various combinations of these three loss functions are tried.
+These loss functions monitor the training process by quantifying the similarity between the ground truth and the predicted masks.
+The Dice loss, related to F1 score, is derived from the Dice coefficient, which calculates the intersecting area between the ground truth and the predicted masks.
+The Dice loss is formulated using (6).
+$l_{Dice}=\frac{FP+FN}{2\times TP+FP+FN}.$ (6)
+In (6), TP stands for the number of true positive pixels, FP for the number of false positive pixels, and FN for the number of false negative pixels.
+The BCE loss minimizes the dissimilarity between the predicted and ground truth masks. The BCE loss can be formulated as (7).
 $\begin{split}l_{BCE}(I,\hat{I})=-\frac{1}{WH}(\sum_{i=0}^{W-1}\sum_{j=0}^{H-1}[I(i,j)log(\hat{I}(i,j))\\ +(1-I(i,j))log(1-\hat{I}(i,j))]).\end{split}$ (7)
-where I represents the ground truth label , $\hat{\textbf{{I}}}$ represents the predicted probability of the foreground.
-Focal loss is used to address class imbalance in segmentation tasks and provide more focus on difficult cases.
-It assigns higher weights to misclassified pixels, while reducing the impact of easy pixels during training.
-The Focal loss is computed as follows
+where I represents the ground truth label , $\hat{\textbf{I}}$ represents the foreground prediction.
+Focal loss is used to concentrate on difficult cases in image segmentation tasks.
+The larger weights are given to the hard part of the image pixels, while the smaller ones are assigned to the easy pixels during training.
+The Focal loss is formulated as follows
 $l_{Focal}=-\frac{1}{N}(\sum_{i=1}^{N}\alpha(1-p_{i})^{\gamma}log(p_{i})).$ (8)
-In Eq.
-9, $\alpha$ is a balancing parameter to control the contribution of each class, and $\gamma$ is the focusing parameter that modulates the rate at which the loss focuses on hard classification cases.
-In training, we use a combination of the Dice loss, BCE loss, and Focal loss as shown in Eq.
-10 to guide the optimization process.
-By minimizing this combined loss during training, our model learns to accurately segment the desired regions of interest.
-$l=l_{Dice}+l_{BCE}+l_{Focal}.$ (10)
+In (8), $\alpha$ is a balancing parameter and $\gamma$ is the focusing parameter.
+In training, we tried using a single loss function and also combinations of these functions, as shown in forms of (9) to monitor the optimization process.
+$l=\alpha l_{Dice}+\beta l_{BCE}+\lambda l_{Focal}.$ (9)
 4 Experiments and Results
 4.1 Dataset Description
-We use the brain anatomy segmentation (Brain US) [24], [27] dataset to evaluate our model.
+We use the brain anatomy segmentation data set (Brain US) [27], [30] to evaluate our model.
 Intraventricular hemorrhage (IVH) resulting in enlargement of the brain ventricles is one of the main causes of preterm brain injury.
@@
 Automatic segmentation of the brain ventricles and septum pellucidum from these US scans is essential for an accurate diagnosis and prognosis of these diseases.
-After obtaining institutional review board (IRB) approval, US scans were taken from 20 different premature neonates (age ¡ 1 year).
+After obtaining institutional review board (IRB) approval, US scans were taken from 20 different premature neonates (age less than 1 year).
 The total number of images collected was 1629 with annotations out of which 1300 were allocated for training and 329 for testing.
@@
 4.2 Evaluation Metrics
-The performance of our segmentation model is evaluated using commonly used metrics: dice score (F1) and intersection over union (IoU) score.
-These metrics provide quantitative measures of the model’s ability to accurately delineate regions of interest.
-Intersection over Union
-IoU is a measure that quantifies the overlap between the predicted binary segmentation mask and the ground truth mask.
-It is calculated as the ratio of the intersection area between the two masks to their union area and is defined as follows
+Tow metrics are used to evaluate the performance of our segmentation model: dice score (F1) and intersection over union score (IoU).
+These metrics provide a quantitative way to evaluate the performance of the segmentation process.
+4.2.1 Intersection over Union
+IoU is a measure that computes the overlap between the predicted segmentation mask and the ground truth mask. It is formulated as follows
 $IoU=\frac{TP}{TP+FP+FN}.$ (10)
-Dice Score
-The Dice score, also referred to as the F1 score, integrates both recall and precision into a single value for evaluation and is defined as follows
+4.2.2 Dice Score
+The Dice score, also referred to as the F1 score, is formulated as follows
 $F_{1}=\frac{2\times TP}{2\times TP+FP+FN}.$ (11)
 4.3 Experimental Setup
-We have developed our segmentation model using Python and have used the TensorFlow and Keras libraries for implementation. To speed up training, we used the high performance NVIDIA RTX 4060Ti 16G.
+We have developed our segmentation model using Python and have used the PyTorch library for implementation. To speed up training, we used the high performance NVIDIA RTX 4060Ti 16G.
 4.4 Hyperparameter Setup
-Our model is trained for 150 epochs for the Brain US dataset, where every epoch represents one complete pass through the entire datasets.
-For the image size, we resize all images to a uniform size of 128×128 pixels, which are input into the model for segmentation.
-During the training phase, we use the Adam optimizer with a learning rate of 0.0001 to optimize the model.
-This choice of optimizer allows us to efficiently update the model’s parameters and enhance convergence during training.
-To ensure a comprehensive assessment, we have divided the data into a split 70% (train), 10% (test) and 20% (validation).
-We have leveraged the training set to optimize the model’s parameters, fine-tune hyperparameters using the validation set, and gauge the model’s ability to generalize to new data by utilizing the testing set.
-4.5 Ablation Study
-Upgrading Model from Scratch
+Our model is trained for 150 epochs for the Brain US dataset.
+For the image size, we resize all images to a uniform size of 128×128 pixels, which are fed into the segmentation model.
+During the training phase, we use the Adam optimizer with a learning rate of 0.00001 to optimize the model.
+To ensure a comprehensive evaluation, we use the common split strategy and divide the data into a split: 70% (train), 10% (test) and 20% (validation).
+All the setups discussed above are usually common in such a training task.
+Figure 3: Results of the ablation study indicate the improvement in model performance with each experimental modification.
+GT and PM are ground truth and predicted mask, respectively.
+Heatmaps demonstrate the improvement of the model’s performance in focusing on the region of interest.The SAL branch filters out the negative impact of query-key matching scores so that the heatmap looks monotonic in the uninformative regions.
+Note that there is no attention map for case i and case ii.
+4.5 Ablation Tests
+4.5.1 Upgrading Model from Scratch
 A series of experiments are conducted to refine our segmentation model and evaluate the impact of upgrading modifications. These experiments include:
@@
   4. iv.
-    Proposed Res-UNet model with the strengths of CBAM and SAL.
+    Proposed Res-UNet model with the strengths of CBAM and HAL.
 The results in Table 2 show the efficacy of each modification in the upgrading way.
-Each model has been trained using the linear combination of Dice, BCE, and Focal loss.
-The CBAM improves performance, while the SAL further enhances accuracy and segmentation quality.
-Both CBAM and SAL increase Dice and IoU by around 3 points with respect to the baseline.
+Each model has been trained using the linear combination of Dice, BCE, and Focal loss with average weights.
+Both CBAM and HAL increase Dice and IoU by around 3 points with respect to the baseline.
 Furthermore, compared to the pure Res-UNet case, the dual attention modules between encoders and decoders contribute about 1.5 to 2 point improvement in terms of both metrics.
-Figure 3: Results of the ablation study indicate the improvement in model performance with each experimental modification.
-GT and PM are Ground Truth and Predicted Mask, respectively.
-Heatmaps demonstrate the improvement of the model’s performance in focusing on the region of interest.The SAL module filters out the negative impact of query-key matching scores so that the heatmap looks monotonic in the uninformative regions.
-Note that there is no attention map for case i and case ii.
-Figure 3 visually illustrates the enhancement achieved with attention mechanisms.
-The combination of CBAM and SAL results in improved performance for both large and small regions of interest, refining feature representations, and capturing both global and local spatial dependencies for accurate segmentation.
-Compared with the CBAM-only case (iii), the full model (iv) performs the segmentation more precisely because the SAL module filters out the ’disturbing’ information from the attention map.
+Fig.
+3 visually illustrates the performance achieved with our combinations of CBAM and HAL attention mechanisms.
+The entire framework can refine feature representations and capture both global and local spatial correlations for accurate segmentation.
+Compared with the CBAM-only case (iii), the full model (iv) performs the segmentation more precisely because the SAL branch filters out the ’disturbing’ information from the attention map.
 Furthermore, the heatmap visualization of the proposed model illustrates the spatial regions where it places more emphasis, showing a close resemblance to the ground truth regions for the Brain US dataset.
-Through these experiments and analysis, we are able to improve our segmentation model iteratively, identifying the most effective modifications and attention mechanisms.
-These advances make a significant contribution to enhancing the accuracy and robustness of our model, positioning it as an advanced solution for segmenting ultrasound images.
-Combinations of Loss Functions
-We have performed a series of experiments utilizing different loss combinations and have consolidated the results in Table 1.
-This table highlights the ablation study we have conducted on the loss functions employed to train our model.
-According to the results in Table 1, we have determined the most efficient loss function during model training and selected the optimal model configuration.
-Our findings reveal that the combination of Dice, BCE, and focal loss yields the highest performance.
+4.5.2 Combinations of Loss Functions
+We have performed some incremental tests on different combinations of loss functions and have consolidated the results in Table 1.
+Based on the results in Table 1, we have compared the Dice and IoU scores for each combination.
+Our conclusion is that the linear combination of all three loss functions produces the best result in our segmentation tests.
+Therefore, we applied this paradigm for all the experiments in the paper.
 Figure 4: Results of the proposed segmentation model on images of the Brain dataset. All the results are obtained by running their respective code.
@@
 The comparative results, which comprehensively evaluate two metrics, are presented in Table 3.
-The models compared in the table are well known in the image segmentation field, such as FCN [7], U-Net [16], UNet++ [17], Res-UNet [25], Axial Attention U-Net [19] and Medical Transformer [4].
+The models compared in the table are well known in the field of image segmentation, such as FCN [7], U-Net [15], UNet++ [16], Res-UNet [28], Axial Attention U-Net [18], and Medical Transformer [4].
 Our proposed method demonstrates superior performance across these models (see Table 3) in terms of Dice score and IoU, indicating better overall segmentation accuracy.
@@
 In addition, the IoU metric, with a value of 81.84, indicates the strong capability of the model to accurately delineate regions of interest.
-Significantly, the proposed method achieves the highest Dice score and IoU out of all the models, suggesting that it excels in terms of segmentation accuracy and overlap with ground truth.
+In conclusion, the proposed model achieves the highest Dice scores out of all the models, suggesting that it performs better than other models.
 Figure 5: Illustration of some of the failure cases of our model. The enclosed regions are the misclassified segmented masks. GT and PM represent the Ground Truth and Predicted Mask, respectively.
@@
 The results are in x(+y) format, where x is the score and y is the improvement over the baseline.
-Metrics (i)U-Net[17] (ii)Res-UNet[27] (iii)Res-UNet (w/CBAM only) (iv)Res-UNet (w/CBAM and SAL(Ours))
+Metrics (i)U-Net[17] (ii)Res-UNet[27] (iii)Res-UNet (w/CBAM only) (iv)Res-UNet (w/CBAM and HAL(Ours))
 Dice 85.37 87.50(+2.13) 87.97(+2.60) 89.04(+3.67)
@@
 IoU 75.02 79.31 79.95 79.61 80.14 81.34 81.84
-Overall, the evaluation results show the superior performance of our proposed model compared to the state-of-the-art models.
-This confirms the effectiveness and robustness of our model in achieving accurate and precise segmentation results, positioning it as a promising solution for various segmentation tasks.
-Figure 4 shows the segmentation results of our proposed model, demonstrating its ability to segment the regions of interest in ultrasound images accurately.
+Fig.
+4 shows the segmentation results of our proposed model, demonstrating its ability to segment the regions of interest in ultrasound images accurately.
 Although the visual result differences between our model and MedT are somehow tiny in most of the masks, it still achieves the best segmentation results by far.
-This indicates that the model focuses its attention on relevant areas, contributing to its accurate segmentation performance.
-Through these comprehensive evaluations and visualizations, our proposed model showcases its potential to significantly improve brain anatomy detection and diagnosis, bringing us closer to the goal of early detection and enhanced patient care.
+This suggests that the model pays attention to only relevant areas, which is beneficial to its segmentation accuracy.
+By these comprehensive evaluations and visualizations, our proposed model demonstrates its ability to significantly improve the detection and diagnosis of brain anatomy , leading to the goal of early alarming.
 4.7 Weakness Analysis
-Our model has demonstrated excellent performance in various image segmentation tasks, as depicted in Table 3 and Figure 4.
+Our model has shown superior performance in various image segmentation tasks, as depicted in Table 3 and Fig.
+4.
 However, it is essential to highlight the failure cases in which non-anatomical regions are misclassified as anatomical ones and vice versa, indicating that the precision and recall are relatively lower.
-It is important to acknowledge the complexity of the dataset used for the evaluation, which presents challenges in achieving perfect segmentation results.
-Figure 5 illustrates the difficult cases which our model encounters, resulting in deviations from the ground truth segmentation.
-These challenges may arise from dataset complexity, variations in image quality, or the presence of ambiguous features that are hard to accurately delineate.
-Despite these challenges, our proposed model demonstrates a valuable contribution to the field of brain anatomy segmentation.
-4.8 Experimentation on the Nuclei Dataset
-To evaluate the effectiveness of our proposed method, we have conducted assessments on another dataset named Nuclei, a well-known collection containing microscopic images of nucleus-stained slides and the corresponding ground truth annotations by expert pathologists.
-The average resolution for Nuclei images and the corresponding ground truth masks is 2000×2000 pixels.
-For the evaluation on this dataset, we have maintained consistency using the same set of hyperparameters that are employed in the evaluation of the Brain US dataset.
-However, the images should be converted to grayscale and resized to the same size as those in the Brain US dataset before being fed to the learning network.
-The segmentation results of the proposed model on the Nuclei dataset are shown in Figure 6.
-The visual results show that the overall performance is worse than those for the Brain US dataset because the segmentation regions are numbers of tiny particle-like areas, which make it hard for the attention to focus.
+Intuitively, the quality of segmentation results is dependent on the complexity of the dataset for training and evaluation.
+Fig.
+5 illustrates the difficulties encountered by our model, leading to deviations from the ground truth.
+These challenges may arise from the complexity of the dataset, variations in image quality, or the presence of ambiguous features that are difficult to delineate accurately.
+Despite these challenges, our proposed model has shown valuable contributions in the field of anatomical segmentation.
 Figure 6: Predicted mask visualization of the proposed model on the Nuclei dataset.
@@
 The enclosed regions are the misclassified segmented masks.
+4.8 Experimentation on the Nuclei Dataset
+In order to evaluate the effectiveness of our proposed method, we performed an evaluation on another dataset named Nuclei, which is a well-known collection that contains bright microscopy images of stained cells and correctly annotated masks by expert pathologists.
+The average resolution of the Nuclei images and their corresponding correct annotation masks is 2000×2000 pixels.
+For the evaluation of this dataset, we used the same set of hyperparameters as used in the evaluation of the Brain US data set to maintain consistency.
+The images should be converted to grayscale and resized to the same size as the images in the Brain US dataset before being fed into the network.
+The segmentation results of the proposed model on the Nuclei dataset are shown in Fig.
+6.
+The visual results show that the overall performance is worse than the performance of the Brain US dataset, as the segmentation area is made up of many tiny grainy regions, which makes it difficult to focus attention.
 5 Conclusion and Future Research
-The current study presents an innovative segmentation technique called the attention-aided Res-UNet model to detect brain anatomies in ultrasound images.
-Our model has delivered satisfactory outcomes, revealing its potential to improve detection and diagnosis.
-Nonetheless, it is crucial to recognize the weakness of our proposed method.
-In the weakness analysis section, several potential areas for future research were identified.
-One particularly promising direction involves investigating more hierarchical architectures that integrate multiple forms of stronger attention modules within the model.
-Such an approach has the potential to improve performance and deepen our understanding of complex brain anatomy detection challenges using complementary insights.
-Another direction for future research is to enhance datasets through techniques such as data enhancements, which can increase their diversity and size, thereby enhancing generalization and robustness.
-In addition, evaluating the robustness of the model could provide valuable insight.
-Performing segmentation on other types of medical images, as we did in Nuclei dataset, would be beneficial in assessing the model’s versatility and applicability in diverse medical imaging tasks.
+The current research proposes an innovative segmentation technique, called attention-based Res-UNet model, to detect brain anatomy in ultrasound images.
+Our model has provided satisfactory results, revealing its potential to improve detection and diagnosis.
+In the weaknesses section, some potential areas for future research were identified.
+One particularly promising way involves studying more advanced architectures that integrate multiple more powerful attention within the model.
+This approach has the potential to improve performance and expand our understanding of the complex challenges in detecting brain anatomy.
+Another direction for future research involves the dataset using techniques such as data augmentation, which can increase their diversity and size, leading to improved generalization and robustness.
+Furthermore, generalizing segmentation on other types of medical images, as we did for the nuclei data set, will assess the versatility and applicability of the model to a variety of medical imaging tasks.
 6 Acknowledgments